OpenStudy (anonymous):

lim sin^2 (x/2) / x^2

OpenStudy (anonymous):

1/4 as x approaches zero.

OpenStudy (anonymous):

What did you have in mind as a approaches something?

OpenStudy (anonymous):

Yes , approaches 0 , sorry . But how did you get thet result?

OpenStudy (anonymous):

That"

OpenStudy (anonymous):

I cheated and used my slave, Mathematica 8.04\[\text{Limit}\left[\frac{\text{Sin}\left[\frac{x}{2}\right]^2}{x^2},x\to 0\right]=\frac{1}{4} \]

OpenStudy (anonymous):

Where did you get that program?

OpenStudy (anonymous):

http://www.wolfram.com/mathematica/how-to-buy/education/students.html Start with the preceding. I paid $250 for the Home Edition. I am not a student or an academic. Special pricing for these folks.

OpenStudy (anonymous):

Okay , thanks . :)

OpenStudy (anonymous):

I might add that the current version 8.04 is a big mother, a 1000 Mb installation file as I recall.

OpenStudy (anonymous):

:P